# LlamAr & GemmAr: Enhancing LLMs Through Arabic Instruction-Tuning

### v1

@@ Abstract
Large language models (LLMs) have greatly impacted the Natural Language Processing (NLP) field, particularly for the English language. These models have demonstrated capabilities in understanding and generating human-like text. The success of language models largely depends on the availability of high-quality instruction datasets, which consist of detailed task descriptions and corresponding responses that are essential for training the models to accurately address a variety of prompts. However, the availability and quality of these resources vary by language. While models perform well in English, they often struggle with languages like Arabic, due to the lack of datasets for fine-tuning Arabic-specific tasks. To address this issue, we introduce InstAr-500k, a new Arabic instruction dataset created by generating and collecting content that covers several domains and instruction types. We then assess this dataset by fine-tuning two open-source models, Llama-3-8B-Instruct and Gemma-7B-IT, on several downstream tasks to scale improvements in their functionality. Based on multiple evaluations, our fine-tuned models achieve state-of-the-art performance on several Arabic NLP benchmarks. These outcomes emphasize the effectiveness of our dataset in elevating the capabilities of language models for Arabic. Our instruction dataset bridges the performance gap between English and Arabic language models by providing resources that amplify Arabic NLP development. Building on this foundation, we developed two state-of-the-art models, LlamAr-8B and GemmAr-7B, which are specifically tuned to excel at a wide range of Arabic NLP tasks.
@@ 1 Introduction
Instruction-tuning has been extensively developed for English, contributing significantly to advancements in language technologies through improved model performance and understanding. Previous studies have highlighted the effectiveness of this approach in improving models’ knowledge and reasoning capabilities, Li et al. [2024]. Despite these achievements, there is a significant disparity in the focus on languages other than English, particularly Arabic Boughorbel & Hawasly [2023]. This gap is important, especially given the digital expansion and the increasing demand for Arabic language technologies.
Our objective is to bridge this gap by developing instruction datasets for Arabic and adapting non-Arabic open-source models, such as Llama3-8B-InstructAI@Meta [2024] and Gemma-7B-IT Team et al. [2024a], for Arabic-specific applications. This initiative aims to improve the accessibility of AI technology for Arabic speakers and contribute to more inclusive technological progress. We initiate our approach by creating an Arabic dataset (InstAr-500k) and then fine-tuning open-source LLM models to function effectively in Arabic. Our methodology involves synthetic data generation, human-crafted data collection, and using the LoRA technique Hu et al. [2021] within the LLaMA Factory framework Zheng et al. [2024] for fine-tuning. The effectiveness of these models is assessed against Arabic NLP benchmarks using evaluation metrics to validate our methods.
By pursuing these efforts, we successfully developed new state-of-the-art models—LlamAr-8B and GemmAr-7B—that demonstrate capabilities in handling a variety of tasks related to the Arabic language. These models were specifically fine-tuned to address the unique syntactic and semantic complexities of Arabic.
Their evaluation scores on various Arabic NLP benchmarks quantitatively reflect the effectiveness of LlamAr-8B and GemmAr-7B. These benchmarks encompass a range of tasks designed to test the models’ understanding of context, nuance, and the intricacies of Arabic grammar. The detailed analysis of these benchmarks, which will be discussed in subsequent sections, provides clear evidence of how LlamAr-8B and GemmAr-7B are set to transform Arabic language processing in diverse applications.
This paper is structured as follows: The Methodology section details the approaches used for dataset construction, including synthetic data generation and the use of the LoRA technique. The Analysis section presents benchmarks and discusses the results of tuning the models on Arabic NLP tasks. The final sections address critical aspects of our research and its broader impact.
@@ 2.1 Instruction-tuning
This makes instruction-tuning a critical step in improving the usability and accessibility of LLMs across different languages and tasks. In our project, we developed Arabic language models by enhancing the performance of non-Arabic open-source models like Llama3-8B-Instruct and Gemma-7B-IT. This was achieved by developing Arabic instruction datasets using both synthetic and real data.
@@ 3.1 Overview
To refine large language models for better performance in Arabic, we relied on a methodology that combines monolingual knowledge distillation Kuulmets et al. [2024] and fine-tuning strategies across a spectrum of datasets—both synthetic and human-crafted to improve the performance of the Llama3-8B-Instruct and Gemma-7B-IT models.
@@ 3.2 Training Data
We offer in this section a detailed explanation of the dataset construction process, emphasizing the used methodologies and providing a quality analysis. The following Table 1 lists datasets used in our study, highlighting their origins and the range of tasks they cover. Each dataset was carefully selected to contribute to the effective training and fine-tuning of our Arabic language models.
@@ 3.2.1 Instruction Dataset Construction
Improvements in the chosen open-source LLMs for Arabic language understanding are centered on the development of an extensive Arabic instruction dataset.
@@
The construction of our dataset is detailed through a pipeline, as illustrated in Figure 1.
This section will explore the various stages of the dataset’s construction, providing both visual and descriptive insights into the different steps involved in preparing the data for the effective training of these models.
@@
    This involved collecting raw text, classification labels, and other pertinent data, which we reformulated to fit instructional needs. Further, we conducted an extensive cleaning and filtering process to remove any irrelevant or low-quality data, ensuring the refined data was suitable for conversion into instruction-response pairs. During this phase, we standardized Arabic characters, numbers, and formatting, and adjusted the dataset based on token count to align with our fine-tuning specifications.
@@ 3.2.2 Quality analysis
    Figure 3 illustrates the token length distributions for both the Llama and Gemma tokenizers, providing insights into the dataset’s characteristics. The distributions highlight the dataset’s extensive scope, with 333,886,144 tokens processed in the inputs and 24,139,403 tokens generated, showcasing the dataset’s productive output. The logarithmic scaling reveals concentrated clusters of outputs and instructions within specific length ranges from 0 to 2,000 tokens, emphasizing the tokenizer’s impact on segmentation in computational linguistics.
    Figure 3: Tokens distribution in instructions and outputs for Gemma and Llama tokenizers.
@@
    The InstAr-500k dataset contains a wide range of categories with a well-balanced distribution across the following areas: Religion, Sports, Politics, Science & Technology, Economy & Finance, Entertainment, History, Health, Geography, and Travel. We used zero-shot classification to assign labels to instructions. We sourced these candidate labels for the zero-shot algorithm from the insights provided in the study referenced in Adelani et al. [2024]. This approach enabled the model to generalize to new, unseen categories based on its acquired knowledge, thus eliminating the time-consuming and subjective process of manual labelling.
@@ 3.3 Fine-tuning
In this subsection, we provide the technique and the parameters employed for Supervised Fine-Tuning (SFT) within the LLaMA Factory framework Zheng et al. [2024]. For this, we used two models: Gemma 7B-IT and Llama3-8B-Instruct. Table 2 details their architectural parameters, including the number of layers, hidden dimensions, and other metrics.
@@
Llama3-8B-Instruct 4096 32 14336 32 8 128 128256 525.336.576 8.030.261.248
Table 2: Models architecture parameters.
In addition to model architecture, the choice of hyperparameters has a prominent role in the fine-tuning process, directly influencing the model’s output. During the fine-tuning of Gemma 7B-IT and Llama3 8B-Instruct models, specific hyperparameters were selected to optimize training. Table 3 summarizes the key hyperparameters, highlighting the calibration necessary for the model training.
@@
In our efforts to boost the performance of Llama3-8B-Instruct and Gemma-7B-IT, we refined quantization techniques by reducing the precision of model parameters. This adjustment enabled the model to prioritize important features during training, resulting in more efficient memory usage and faster training times.
@@ 3.4.2 Fine Tuning environment
    We used the Open Arabic LLM Leaderboard (OALL)Elfilali et al. [2024] to track progress and rank our models’ performance. LLMs on the OALL are evaluated with LightEval, a unified framework from the Hugging Face, to test and assess causal language models across multiple evaluation tasks. This includes Arabic translations of benchmarks like MMLU Hendrycks et al. [2020], Exam Hardalov et al. [2020], ARC-Challenge Clark et al. [2018], ARC-Easy Clark et al. [2018], BOOLQ Clark et al. [2019], COPA Roemmele et al. [2011], HellaSwag Zellers et al. [2019], OPENBOOK-QA Mihaylov et al. [2018], PIQA Bisk et al. [2020], RACE Lai et al. [2017], SCIQ Welbl et al. [2017], and TOXIGEN Hartvigsen et al. [2022]. The leaderboard also features benchmarks specifically created for Arabic and its cultural context, such as AlGhafa Almazrouei et al. [2023] and ACVA Huang et al. [2023].
@@ 4.1 Benchmarks
Our study applied a set of benchmarks to evaluate the performance of our models across multiple domains. These benchmarks were selected to cover a broad range of tasks, ensuring a diverse assessment. Key components of our evaluation included the Arabic MMLU Benchmark, which is the translated version of MMLU (Massive Multi-task Language Understanding) provided by the OALL team [^4] https://huggingface.co/OALL. This benchmark is a standardized method for assessing AI performance on a range of tasks, from simple mathematics to complex legal reasoning in Arabic. It consists of 57 tasks across numerous domains, including elementary mathematics, history, computer science, and law, requiring models to demonstrate a broad knowledge base and problem-solving skills.
@@ 4.2 Results
The results from various models on the Open Arabic LLM Leaderboard (OALL) [^6] https://huggingface.co/spaces/OALL/Open-Arabic-LLM-Leaderboard reveal diverse performances across multiple Arabic NLU tasks. Our fine-tuned models, LlamAr-8B and GemmAr-7B, exhibit strong performances with average scores of 62.41% and 60.92%, respectively.
Figure 5: Performance scores comparison of GemmAr-7B, LlamAr-8B, Llama-3-8B-Instruct, and Gemma-7B across different benchmarks.
Figure 5 illustrates the performance of several key benchmarks, including Arabic_MMLU, AlGhafa, ACVA, RACE_Ar, and COPA_Ar. These benchmarks are compared against some of the top-performing models on the OALL leaderboard, such as Llama3-8B-Instruct AI@Meta [2024], openchat-3.5 Wang et al. [2024], Qwen2-7B-Instruct [^7] https://huggingface.co/Qwen/Qwen2-7B-Instruct and Gemma-7B Team et al. [2024b]. The results highlight that our models, LlamAr-8B and GemmAr-7B, stand out as the best-performing models, with higher scores for Arabic tasks.
Our models demonstrate a strong performance in the Arabic_MMLU, AlGhafa, and ACVA benchmarks, which evaluate multi-tasking abilities, contextual understanding, and vocabulary. They also outperform the other models in most benchmarks, particularly in the COPA_Ar and RACE_Ar, which assess commonsense reasoning and reading comprehension respectively.
Benchmark scores based on OALL evalset
Benchmarks GemmAr-7b (ours) LlamAr-8b(ours) suzume-llama-3-8B [^8] https://huggingface.co/lightblue/suzume-llama-3-8B-multilingual-orpo-borda-top25 Gemma-7B [^9] https://huggingface.co/google/gemma-7b openchat-3.5101010https://huggingface.co/openchat/openchat_3.5 Llama-3-8B-Instruct [^11] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct Qwen2-7B-Instruct [^12] https://huggingface.co/Qwen/Qwen2-72B-Instruct
Arabic_MMLU 46.82 42.58 42.47 44.22 45.36 41.25 46.73
ACVA 75.86 79.53 39.83 40.22 44.19 39.87 39.55
AlGhafa 71.76 64.47 45.88 47.88 48.85 46.79 49.68
Arabic_Exams 43.76 41.71 43.76 41.89 43.57 43.94 47.85
ARC_challenge_okapi _Ar 67.84 49.74 38.01 42.15 42.24 39.22 47.06
ARC_easy_Ar 79.73 59.17 38.45 45.60 43.86 39.25 46.95
Boolq_Ar 81.56 64.66 62.11 67.94 81.41 62.11 77.88
Copa_ext_Ar 75.55 66.66 61.11 47.77 50.00 52.22 54.44
Hellaswag_okapi_Ar 30.33 27.92 28.79 26.90 31.65 29.08 32.89
OpenBook_QA_ext_Ar 60.80 50.30 42.62 41.41 51.11 45.45 49.09
Piqa_Ar 65.24 61.04 61.81 46.46 64.53 60.17 66.77
Race_Ar 63.40 51.41 38.40 38.36 43.88 38.38 42.44
Sciq_Ar 68.84 86.63 53.16 41.00 42.01 46.13 41.55
Toxigen_Ar 74.11 45.02 43.20 56.14 67.48 43.20 69.19
Total Average score normalized 62.41 60.92 41.89 42.94 45.56 41.36 44.44
Table 5: Benchmarks scores based on OALL evalset.
@@ 5 Related Work
Fine-tuning large language models presents several challenges, such as potential knowledge erosion, where modifying all parameters can lead to forgetting previously learned tasks Kemker et al. [2018]. To address this issue, Han et al. [2024] introduced Parameter Efficient Fine-Tuning (PEFT), which modifies only a subset of parameters to help retain previously acquired knowledge more effectively Pu et al. [2023]. One such PEFT technique, Low-Rank Adaptation of Large Language Models (LoRA), proposed by Hu et al. [2021], reduces the number of trainable parameters, optimizing the fine-tuning process. By applying LoRA to pre-trained models, high performance can be achieved with minimal computational costs.
@@ 6 Conclusion
In our continuous pursuit of improving Arabic NLP, we are delighted to introduce LlamAr-8B and GemmAr-7B, two state-of-the-art fine-tuned versions of Large Language Models specifically developed for Arabic using our crafted Arabic-instructed dataset InstAr-500k. Through evaluations, LlamAr-8B and GemmAr-7B showcased their abilities, securing top positions among leading models on the OALL leaderboard. By combining the strength of LLMs with our customized training approach, we aim to remove any barrier that hinders the performance of Arabic language tasks and make LlamAr-8B and GemmAr-7B trusted companions for researchers and Arabic LLM enthusiasts. We are excited about the potential impact of these models and look forward to witnessing the different use cases that will be built on LlamAr-8B and GemmAr-7B.

### v2

# GemmAr: Enhancing LLMs Through Arabic Instruction-Tuning

@@ Abstract
Large language models (LLMs) have greatly impacted the natural language processing (NLP) field, particularly for the English language. These models have demonstrated capabilities in understanding and generating human-like text. The success of language models largely depends on the availability of high-quality instruction datasets, which consist of detailed task descriptions and corresponding responses that are essential for training the models to address a variety of prompts accurately. However, the availability and quality of these resources vary by language. While models perform well in English, they often need help with languages like Arabic, due to the lack of datasets for fine-tuning Arabic-specific tasks. To address this issue, we introduce InstAr-500k, a new Arabic instruction dataset created by generating and collecting content that covers several domains and instruction types. We assess this dataset by fine-tuning an open-source Gemma-7B model on several downstream tasks to improve its functionality. Based on multiple evaluations, our fine-tuned model achieves excellent performance on several Arabic NLP benchmarks. These outcomes emphasize the effectiveness of our dataset in elevating the capabilities of language models for Arabic. Our instruction dataset bridges the performance gap between English and Arabic language models by providing resources that amplify Arabic NLP development. Building on this foundation, we developed a model, GemmAr-7B-V1, specifically tuned to excel at a wide range of Arabic NLP tasks.
@@ 1 Introduction
Instruction-tuning has been extensively developed for English, contributing significantly to advancements in language technologies through improved model performance and understanding. Previous studies have highlighted the effectiveness of this approach in improving models’ knowledge and reasoning capabilities Li et al. [2024]. Despite these achievements, there is a significant disparity in the focus on languages other than English, particularly Arabic Boughorbel & Hawasly [2023]. This gap is important, especially given the digital expansion and the increasing demand for Arabic language technologies.
We aim to bridge this gap by developing instruction datasets for Arabic and adapting non-Arabic open-source models, such as Gemma-7B-IT Team et al. [2024], for Arabic-specific applications. This initiative aims to improve the accessibility of AI technology for Arabic speakers and contribute to more inclusive technological progress. We initiate our approach by creating an Arabic dataset (InstAr-500k) and then fine-tuning open-source LLM models to function effectively in Arabic. Our methodology involves synthetic data generation, human-crafted data collection, and using the LoRA technique Hu et al. [2021] within the LLaMAFactory framework Zheng et al. [2024] for fine-tuning. The effectiveness of these models is assessed against Arabic NLP benchmarks using evaluation metrics to validate our methods.
By pursuing these efforts, we successfully developed a new model GemmAr-7B-V1 that demonstrates capabilities in handling a variety of tasks related to the Arabic language. This model was specifically fine-tuned to address the unique syntactic and semantic complexities of Arabic.
Its evaluation scores on various Arabic NLP benchmarks quantitatively reflect the effectiveness of GemmAr-7B-V1. These benchmarks encompass a range of tasks designed to test the models’ understanding of context, nuance, and the intricacies of Arabic grammar. The detailed analysis of these benchmarks, which will be discussed in subsequent sections, provides clear evidence of how GemmAr-7B-V1 is set to transform Arabic language processing in diverse applications.
This paper is structured as follows: The Methodology section details the approaches used for dataset construction, including synthetic data generation and the use of the LoRA technique. The Analysis section presents benchmarks and discusses the results of tuning the model on Arabic NLP tasks. The final sections address critical aspects of our research and its broader impact.
@@ 2.1 Instruction-tuning
This makes instruction-tuning a critical step in improving the usability and accessibility of LLMs across different languages and tasks. In our project, we developed an Arabic language model by enhancing the performance of non-Arabic open-source models like Gemma-7B-IT. This was achieved by developing an Arabic instruction dataset InstAr-500k using both synthetic and real data.
@@ 3.1 Overview
To refine large language models for better performance in Arabic, we relied on a methodology that combines monolingual knowledge distillation Kuulmets et al. [2024] and fine-tuning strategies across a spectrum of datasets—both synthetic and human-crafted to improve the performance of the Gemma-7B-IT model.
@@ 3.2 Training Data
We offer in this section a detailed explanation of the dataset construction process, emphasizing the used methodologies and providing a quality analysis. The following Table 1 lists datasets used in our study, highlighting their origins and the range of tasks they cover. Each dataset was carefully selected to contribute to the effective training and fine-tuning of our Arabic language model.
@@ 3.2.1 Instruction Dataset Construction
Improvements in the chosen open-source LLMs for Arabic language understanding are centered on developing an extensive Arabic instruction dataset. The construction of our dataset is detailed through a pipeline, as illustrated in Figure 1.
@@
This section will explore the various stages of the dataset’s construction, providing visual and descriptive insights into the different steps involved in preparing the data for the effective training of these models.
@@
    This involved collecting raw text, classification labels, and other pertinent data, which we reformulated to fit instructional needs.
    Further, we conducted an extensive cleaning and filtering process to remove any irrelevant or low-quality data, ensuring the refined data was suitable for conversion into instruction-response pairs. During this phase, we standardized Arabic characters, numbers, and formatting, and adjusted the dataset based on token count to align with our fine-tuning specifications.
@@ 3.2.2 Quality analysis
    Figure 3: Tokens distribution in instructions and outputs for Gemma tokenizer.
    Figure 3 illustrates the token length distributions for Gemma tokenizer, providing insights into the dataset’s characteristics. The distributions highlight the dataset’s extensive scope, with 333,886,144 tokens processed in the inputs and 24,139,403 tokens generated, showcasing the dataset’s productive output. The logarithmic scaling reveals concentrated clusters of outputs and instructions within specific length ranges from 0 to 2,000 tokens, emphasizing the tokenizer’s impact on segmentation in computational linguistics.
@@
    The InstAr-500k dataset contains a wide range of categories with a well-balanced distribution across the following areas: Religion, Sports, Politics, Science & Technology, Economy & Finance, Entertainment, History, Health, Geography, and Travel. We used zero-shot classification to assign labels to instructions. We sourced these candidate labels for the zero-shot algorithm from the insights provided in the study referenced in Adelani et al. [2024]. This approach enabled the model to generalize to new, unseen categories based on its acquired knowledge, thus eliminating the time-consuming and subjective process of manual labeling.
@@ 3.3 Fine-tuning
In this subsection, we provide the technique and the parameters employed for supervised fine-tuning (SFT) within the LLaMAFactory framework Zheng et al. [2024]. For this, we used the Gemma 7B-IT model. Table 2 details its architectural parameters, including the number of layers, hidden dimensions, and other metrics.
@@
Table 2: Model architecture parameters.
In addition to model architecture, the choice of hyperparameters has a prominent role in the fine-tuning process, directly influencing the model’s output. During the fine-tuning of the Gemma 7B-IT model, specific hyperparameters were selected to optimize training. Table 3 summarizes the key hyperparameters, highlighting the calibration necessary for the model training.
@@
In our efforts to boost the performance of Gemma-7B-IT, we refined quantization techniques by reducing the precision of model parameters. This adjustment enabled the model to prioritize important features during training, resulting in more efficient memory usage and faster training times.
@@ 3.4.2 Fine Tuning environment
    We tracked progress and ranked our model’s performance using the Open Arabic LLM Evaluation Leaderboard (OALL) Elfilali et al. [2024]. LLMs on the OALL are evaluated with LightEval, a unified framework from the Hugging Face, to test and assess causal language models across multiple evaluation tasks. This includes Arabic translations of benchmarks like MMLU Hendrycks et al. [2020], Exam Hardalov et al. [2020], ARC-Challenge Clark et al. [2018], ARC-Easy Clark et al. [2018], BOOLQ Clark et al. [2019], COPA Roemmele et al. [2011], HellaSwag Zellers et al. [2019], OPENBOOK-QA Mihaylov et al. [2018], PIQA Bisk et al. [2020], RACE Lai et al. [2017], SCIQ Welbl et al. [2017], and TOXIGEN Hartvigsen et al. [2022]. The leaderboard also features benchmarks specifically created for Arabic and its cultural context, such as AlGhafa Almazrouei et al. [2023] and ACVA Huang et al. [2023].
@@ 4.1 Benchmarks
Our study applied a set of benchmarks to evaluate the performance of our model across multiple domains. These benchmarks were selected to cover a broad range of tasks, ensuring a diverse assessment. Key components of our evaluation included the Arabic MMLU Benchmark, which is the translated version of MMLU (Massive Multi-task Language Understanding) provided by the OALL team [^4] https://huggingface.co/OALL. This benchmark is a standardized method for assessing AI performance on a range of tasks, from simple mathematics to complex legal reasoning in Arabic. It consists of 57 tasks across numerous domains, including elementary mathematics, history, computer science, and law, requiring models to demonstrate a broad knowledge base and problem-solving skills.
@@ 4.2 Results
The results from various models on the Open Arabic LLM Leaderboard (OALL) [^6] https://huggingface.co/spaces/OALL/Open-Arabic-LLM-Leaderboard reveal diverse performances across multiple Arabic Natural Language Understanding (NLU) tasks. Our fine-tuned model, GemmAr-7B-V1, exhibits strong performances with average scores of 47.27%.
Figure 5: Performance Scores Comparison of GemmAr-7B-V1, AceGPT-7B-chat, Gemma 7B-IT across different Benchmarks.
Figure 5 illustrates the performance of several key benchmarks, including Arabic_MMLU, AlGhafa, EXAMS, and BOOLQ. These benchmarks are compared against the base model, Gemma 7B-IT Team et al. [2024], and AceGPT 7B-chat Huang et al. [2023] on the OALL leaderboard. The results in Table 5, highlight that our model, GemmAr-7B-V1, stands out as the best-performing model, with higher scores for Arabic tasks. It demonstrates strong performance in the Arabic_MMLU, AlGhafa, and Toxigen_Ar benchmarks, which evaluate multi-tasking abilities, contextual understanding, and vocabulary. It also outperforms the other models in most benchmarks, particularly in the SCIQ_Ar and RACE_Ar, which assess commonsense reasoning and reading comprehension respectively.
Benchmark Scores per model
Benchmarks GemmAr-7B-V1 (ours) Gemma7B-IT AceGPT 7B-chat
Arabic_MMLU 44.45 35.27 33.94
ACVA 42.76 40.29 46.84
AlGhafa 47.56 47.27 45.29
Arabic_Exams 45.81 31.47 35.38
ARC_Challenge 38.36 34.05 32.16
ARC_Easy 38.54 35.07 34.39
Boolq_Ar 80.40 59.69 66.90
Copa_ext_Ar 48.88 51.11 55.56
Hellaswag_Ar 26.65 26.42 26.54
OpenBook_QA 40.61 41.41 38.59
Piqa_Ar 57.06 54.88 54.45
Race_Ar 35.46 35.20 33.15
Sciq_Ar 50.35 40.80 44.72
Toxigen_Ar 64.81 57.22 43.10
Total Average score normalized 47.27 42.15 42.21
Table 5: OALL evalset 0-shot benchmark scores.
@@ 5 Related Work
Fine-tuning large language models presents several challenges, such as potential knowledge erosion, where modifying all parameters can lead to forgetting previously learned tasks Kemker et al. [2018]. To address this issue, Han et al. [2024] introduced Parameter Efficient Fine-Tuning (PEFT), which edits only a subset of parameters to help retain previously acquired knowledge more effectively Pu et al. [2023]. One such PEFT technique, Low-Rank Adaptation of Large Language Models (LoRA), proposed by Hu et al. [2021], reduces the number of trainable parameters, optimizing the fine-tuning process. By applying LoRA to pre-trained models, high performance can be achieved with minimal computational costs.
@@ 6 Conclusion
In our continuous pursuit of improving Arabic NLP, we are delighted to introduce GemmAr-7B-V1, a fine-tuned version of a Large Language Model specifically developed for Arabic using our crafted Arabic-instructed dataset InstAr-500k. By combining the strength of the LLM with our customized training approach, we aim to remove any barrier that hinders the performance of Arabic language tasks and make GemmAr-7B-V1 trusted companions for researchers and Arabic LLM enthusiasts. We are excited about the potential impact of this model and look forward to witnessing the different use cases that will be built on GemmAr-7B-V1.